OpenStudy (amonoconnor):

I need help in evaluating the integral of the following: t(1 + t)^(1/2)dt, from x = 0 to 3. Can one distribute the first "t" of the integral through the parenthesis? I tried leaving it as is, and evaluating through u-substitution, with u = (1 + t), and had a "t" I couldn't get rid of :/ Any and all help is greatly appreciated! :)

OpenStudy (abb0t):

find a substitution for "\(t\)"

OpenStudy (amonoconnor):

Here is what I have worked through thus far: u = (1 + t) du = ((0) + t^(1-1)) = (t^0) = (1)dt = dt Making the integral: (u^(1/2)) * t * du Right?

OpenStudy (amonoconnor):

If you want a good laugh... Read the bio on the smart score page for abb0t lol.

OpenStudy (wolfe8):

._. How can you integrate t with respect to x? Wouldn't the t's just be constants then? Am I missing something?

OpenStudy (amonoconnor):

Sorry, I completely missed that... I'm so used to typing dx, it slipped my mind- good catch!

OpenStudy (wolfe8):

Ah ok. Well since you have u=t+1, you will replace that last t with u-1

OpenStudy (amonoconnor):

Ohh! Tell me if this works!! du = dt So does... du(t) = du(t) ?

OpenStudy (amonoconnor):

Oh, okay.. I see how that works. :) does my way work too?

OpenStudy (amonoconnor):

Oh wait... That still leaves a t (my way). So... Combination of both thoughts: du(t) = du(u - 1)

OpenStudy (wolfe8):

I'm not sure what you mean with du(t) = du(t). You just substitute all the t into u terms and then integrate term by term.

OpenStudy (wolfe8):

Here: You have u= 1+t so t= u-1 and du = dt. So substitute them you get integral (u-1) sqrt(u) du

OpenStudy (amonoconnor):

Sweet, that's what I was putting together, but was poorly articulating. :) thanks you!

OpenStudy (wolfe8):

Ah ok. So you should be able to solve it pretty easily now.

OpenStudy (amonoconnor):

Should I distribute the u^(1/2) through the (u - 1)? Can I do that?

OpenStudy (amonoconnor):

(u^(3/2) - u^(1/2))du ? Can I do that, then take the anti-derivative?

OpenStudy (wolfe8):

Yup yup.

OpenStudy (amonoconnor):

*working problem...

OpenStudy (abb0t):

@nincompoop

OpenStudy (amonoconnor):

Do you disagree with something abbot ?

OpenStudy (amonoconnor):

I got, as a final answer, 182/15, or 12.133333. Is this what you got?

OpenStudy (wolfe8):

Hmm I got 116/15

OpenStudy (wolfe8):

After integrating and substituting I got 2/15 (t+1)^(3/2) (3 t-2)

OpenStudy (amonoconnor):

Dang. Okay. Umm... So, after that u^(1/2) was distributed the integral was... (u^(3/2) - u^(1/2))du Anti-deriviative taken... = u^(5/2) - u^(3/2) = 2/5(u^(5/2)) - 2/3(u^(3/2)) = 2/5(1 + t)^(5/2) - 2/3(1 + t)^(3/2) = 2/5(1 + 3)^(5/2) - 2/3(1 + 0)^(3/2) = 2/5(4)^(5/2) - 2/3(1)^(3/2) = 2/5(32) - 2/3(1) = (64/5) - (2/3) = (192/15) - (10/15) = 182/15 = 12.13333 What did I do wrong?

OpenStudy (amonoconnor):

Sorry that took so long to type:/

OpenStudy (wolfe8):

It's ok. When you put in the limits, you put in the upper limit in the whole antiderivative then minus the lower limit in the whole antiderivative. I see you doing half and half here. Get what I mean?

OpenStudy (wolfe8):

So your final antiderivative is 2/5(1 + t)^(5/2) - 2/3(1 + t)^(3/2) First you put t=3 in the whole equation. Then t=0 in the whole equation. And subtract the first with the second.

OpenStudy (amonoconnor):

Yes, I totally see what you mean:) * stupid Amon... Thanks again:)

OpenStudy (amonoconnor):

I don't understand how I make such stupid mistakes, on such simple problems!!

OpenStudy (wolfe8):

Nah it's fine this is what learning is. Don't beat yourself up